OpenStudy (paigelovestaylor):

An upward force acts on a proton as it moves with a speed of 2.0 × 105 meters/second through a magnetic field of 8.5 × 10-2 tesla running from east to west. What is the force and direction of the force on the proton?

OpenStudy (shamim):

can u draw it?

OpenStudy (paigelovestaylor):

no?

OpenStudy (osprey):

Here's my guess. Bqv is the force on a charged particle moving in a magnetic field. The direction is either by FLEMING'S LEFT HAND MOTOR RULE, or by vector cross products. (8.5 × 10-2)x(1.6x10^-19)X(2x10^5)=0.0000000000000272N= 2.72^10^-14N The force is at right angles to both the field direction and to the direction of the velocity. There's a couple of slight problems, though. One is that protons are positively charged so they'd be deflected in the opposite direction to the direction electrons would be deflected. So, your direction is either N to S or it's S to N. Or looking at the q again, it DOES say UPWARD. So, I guess that the field is East-West, the protons are going either N-S or S-N, whichever direction gives the UPWARD bit in the q. Another problem is with the numbers. The answer here seems a bit low ... but that's what I think the numbers say. And this assumes low speeds, and is a low speed (believe it or not) calculation, that is it is non relativistic. Hope I'm rightish on this one. You can always tell physics students in an exam when they hit elec and mag by the contortions they put their hands in to try to work out Fleming's motor and dynamo rules. http://perendis.webs.com

OpenStudy (icrazyaliens):

To find the Force use this formula \[F = qvB\] \[q = (+1.6 \times 10^-19) C\] Right Hand Rule 1. Index finger shows you the velocity which the direction of moving 2. Middle finger shows you the direction of the magnetic field 3. Thumb finger shows you the direction of the force. Now, the problem said that the proton moves from East to West, by using the right hand rule, you will find the force will be upward. Note: It is a proton so it will deflected upward and the force is positive , if it is an Electron it will be attracted downward.

OpenStudy (osprey):

@IcrazyAliens Interesting take ... differs from mine ... I thought it was the left hand motor rule. oh well, maybe the numbers are right ... @PaigeLovesTaylor

OpenStudy (irishboy123):

IMHO: drawing it is the best idea +1 shamim or one can assume that it is all orthogonal, and into or out of the page (in which case one can just roll out equations and invent answers) but even then, the Lorentz force law says that: \(\vec F = q (\vec v \times \vec B)\) which contains all of the above info in a very simple format :-)

OpenStudy (osprey):

I tend to remember my version of this lot as "Bil Bev and BAnI", throw my fingers around and hope for the best. very scientific, I know.